### v2

# Towards Efficient and Scalable Implementation of Differentially Private Deep Learning

Abstract
Differentially private stochastic gradient descent (DP-SGD) is the standard algorithm for training machine learning models under differential privacy (DP). The most common DP-SGD privacy accountants rely on Poisson subsampling for ensuring the theoretical DP guarantees. Implementing computationally efficient DP-SGD with Poisson subsampling is not trivial, which leads to many implementations ignoring this requirement. We conduct a comprehensive empirical study to quantify the computational cost of training deep learning models under DP given the requirement of Poisson subsampling, by re-implementing efficient methods using Poisson subsampling and benchmarking them. We find that using the naive implementation DP-SGD with Opacus in PyTorch has between 2.6 and 8 times lower throughput of processed training examples per second than SGD. However, efficient gradient clipping implementations with e.g. Ghost Clipping can roughly halve this cost. We propose alternative computationally efficient ways of implementing DP-SGD with JAX that are using Poisson subsampling and achieve only around 1.2 times lower throughput than SGD based on PyTorch. We highlight important implementation considerations with JAX. Finally, we study the scaling behaviour using up to 80 GPUs and find that DP-SGD scales better than SGD. We share our re-implementations using Poisson subsampling at https://github.com/DPBayes/Towards-Efficient-Scalable-Training-DP-DL.
1 Introduction
Training data of machine learning (ML) models can be vulnerable to extraction (Balle et al., 2022; Carlini et al., 2021). Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for formalizing the privacy leakage of training data in ML and mitigating the risk of privacy attacks on the training data. DP is deployed in many applications involving sensitive data (Abowd, 2018; Cormode et al., 2018).
The established algorithm for integrating DP into the training pipeline of deep learning models is DP stochastic gradient descent (DP-SGD) (Rajkumar & Agarwal, 2012; Song et al., 2013; Abadi et al., 2016), which is the DP adaptation of SGD (see also Algorithm 1). DP-SGD has two major drawbacks in comparison to SGD: higher computational cost and loss in utility. DP-SGD requires more memory and is computationally more expensive due to the per-example clipping. The utility in comparison to non-DP training drops but this can be mitigated to certain extend by using larger batch sizes (Räisä et al., 2024) and training longer (Ponomareva et al., 2023) which further increases the computational cost.
Standard DP bounds assume so-called Poisson subsampling, where each example is selected at each iteration independently with a fixed probability. This implies that different minibatches will be of different size, and makes efficient implementation more difficult. As a result, many existing implementations forego proper implementation of Poisson subsampling. Recent research (Lebeda et al., 2024) shows that such implementations may have significantly weaker privacy guarantees than claimed under Poisson subsampling assumption.
List of contributions In this work we perform an extensive empirical study on the computational efficiency of DP-SGD. We will focus on fine-tuning a wide range of large computer vision classification models but our findings can be applied any other large models that are trained or fine-tuned with DP-SGD. Our main contributions are the following:
1.
We re-implement all DP-SGD methods with Poisson subsampling that is fully DP and share the source code at https://github.com/DPBayes/Towards-Efficient-Scalable-Training-DP-DL
2.
We find that non-optimized training with DP-SGD costs betweem 2.6 and 3.2 times more than non-private training for ViT and 4 to 8 times for ResNets (See Section 4). We identify the reasons that lead to the higher computational cost of DP-SGD using profiling.
3.
We benchmark different strategies that can reduce this cost drastically up to a level that matches the non-optimized non-private training (See Figure 1 for an overview): 1 More efficient gradient clipping implementations of DP-SGD (See Section 5.1). 2 Lower Precision with TF32 (See Section 5.2).
4.
We propose a JAX implementation relying on proper Poisson sampling that is not prone to re-compilation and outperforms a naive implementation in terms of throughput (See Section 6).
5.
We scale up the training to 80 GPUs and find that DP-SGD scales better than non-private training (See Section 7).
2 Background
This section will explain the main DP-SGD algorithm and optimizations to alleviate its computational cost.
2.1 DP-SGD Algorithm
Algorithm 1 is the original DP-SGD algorithm, with virtual batching, as proposed by Abadi et al. (2016).
Virtual Batching distinguishes between logical and physical batches. Logical batches are divided into multiple physical batches to allow taking optimizer steps with many samples without running out of memory. In our experiments, we typically sample logical batch sizes of size while the memory only fits samples at a time. Implementing DP-SGD with virtual batching Algorithm 1 does not modify the privacy accounting. The amount of added noise is the same and does not affect the model utility (Ponomareva et al., 2023).
Opacus Our baseline uses the PyTorch (Ansel et al., 2024) library Opacus (Yousefpour et al., 2021). Opacus is the most mature DP-SGD framework out of all considered implementations. It supports nearly all neural layers that are compatible with DP training. Opacus implements per-sample clipping without any additional optimizations.
Opacus implements virtual batching algorithm in their BatchMemoryManager. The privacy engine of Opacus will sample the logical batches, as in the original DP-SGD algorithm, and then divide them into physical batches.
The other implementations considered in our experiments do not support virtual batching out-of-the-box.
Poisson subsampling Interestingly, Bu et al. (2022) and Bu et al. (2023) never mention Poisson subsampling in their works of Mix Ghost clipping and Book Keeping. Even more, Bu et al. (2022) state that it has a speed-up of times against other algorithms with a fixed batch size, which would affect the privacy accounting method. The same happens in practice for JAX implementations (De et al., 2022), where the sampling is done by shuffling the dataset and using each sample once per epoch. While this makes efficient implementation easier, it does not use the correct Poisson subsampling assumed by the privacy accounting methods, and therefore the implementation might have significantly weaker privacy properties than claimed (Lebeda et al., 2024).
To make a fair comparison between all methods, we implement the Poisson subsampling, the same way Opacus does it, for all frameworks and a new custom Batch Memory Manager to flag when it is time to take a step. This way, all the experiments are seeded and have the same logical batch sizes.
2.2 DP-SGD Gradient Clipping Optimizations
We benchmark five types of clipping methods. Table A1 shows which clipping optimizations we are benchmark against the library or framework that implements it.
Ghost clipping computes the loss gradient norm after the backpropagation optimization and then reweights the loss to update the clipped gradients. Its main contribution is memory saving at the cost of adding another backward pass (Li et al., 2022).
Mixed Ghost clipping (Bu et al., 2022) builds on-top of Ghost clipping. It implements the ghost clipping technique for convolutional layers. Its main contribution is that the algorithm will decide when to clip the gradients using per-example or ghost. This difference matters because the ghost clipping is less efficient when the layer’s input dimensions are too big. E.g., for ResNets, each clipping method will be applied for half of the layers. The first layers will be clipped using the the per-example and then ghost clipping in the bottom layers. As the model goes deeper, the feature size decreases, and the number of channels increases, prioritizing ghost clipping (Bu et al., 2022).
Book Keeping (Bu et al., 2023) uses all the previous techniques but requires only one backpropagation pass without explicitly calculating the per-example gradients. It avoids the second pass that ghost clipping does by reusing the intermediate results of the output gradients to calculate the sum of the clipped gradients and the clipping factor. Book Keeping can also be implemented together with the mix optimization. It also implements another technique called MixOpt, which does the same evaluation as the mix ghost clipping, but also determines whether doing a second backward pass is more efficient.
Implementations of optimized DP-SGD PrivateVision (Bu et al., 2022), and FastDP (Bu et al., 2023) are PyTorch-based implementations. While both PrivateVision and FastDP implement ghost clipping and its variants, all FastDP implementations use the Book Keeping method. By the time of the submission, Opacus ghost clipping implementation was not ready for production. We tested it but excluded it from the analysis.
JAX We compare all implementations with two JAX implementations: The first one is a native JAX (Bradbury et al., 2018) implementation that clips the per-sample gradients with Optax (DeepMind et al., 2020) without utilizing any further optimization. This naive implementation in JAX is prone to recompiling due to changing tensor sizes caused by the Poisson subsampling. The second implementation which we call Masked DP-SGD overcomes this issue by processing always full physical batches and then masking gradients of samples that are not required (see Algorithm 2). This implementation computes always slightly more gradients but is not prone to repeated recompilation.
3 Experiment Overview
Metrics We compare the throughput, defined as how many samples can be processed per second during training, and the maximum physical batch size, reached before running out of memory.
Dataset We benchmark with the CIFAR100 (Krizhevsky & Hinton, 2009) resized to 224x224.
Models We train two families of models: Vision Transformer (ViT) (Dosovitskiy et al., 2021) and ResNet (Kolesnikov et al., 2020) (See Table 1). Both are pre-trained on ImageNet-21k (Russakovsky et al., 2015).
| Vision Transformer (ViT) | ResNet | ||
| Type | # of Parameters | Type | # of Parameters |
| Tiny | 5.7 M | 23.7 M | |
| Small | 22.1 M | 42.7 M | |
| Base | 86.6 M | 211.8 M | |
| Large | 304.3 M | 382.4 M | |
| Huge | 630.8 M | 929.2 M | |
Parameterization While parameter-efficient fine-tuning of some parts of the model has been shown to be effective under DP (Tobaben et al., 2023; Yu et al., 2022), our work focuses on the computational efficiency of DP-SGD and thus we consider the worst-case scenario of fine-tuning all parameters of the model. Furthermore, any training from scratch requires training all parameters.
Hyperparameters We train for four optimizer steps with a sampling rate of (expected batch size of 25000), which allows us to test the experiments quickly with a realistic high batch size (Ponomareva et al., 2023; Räisä et al., 2024). We do not focus on finding the best possible utility, which requires training for many more epochs (See Table A3 for the accuracy after training for four steps).
Environment specifications We use two GPU architectures: NVIDIA V100 (32 GB VRAM) and A100 (40 GB VRAM) with identical Python environments. Each node contains four GPUs. We use 16 CPU workers for data loading. In the distributed case of more than one GPU, we cannot use multiple workers.
Source code We provide the code for reproducing the experiments at https://github.com/DPBayes/Towards-Efficient-Scalable-Training-DP-DL.
4 What is the cost of DP in Deep Learning?
In this section we will quantify the computational cost of deploying DP training. We do this by comparing the throughputs and maximum physical batch sizes between the non-private training with PyTorch and private training with Opacus, which is the most used DP-SGD implementation. Additionally, we identify the reasons for the higher computational cost of DP-SGD through profiling.
4.1 Throughput and Maximum Batch Size Comparison
We compare relative throughput (Figure 2) and the maximum physical batch size (Figure 3) between DP-SGD (Opacus) and non-private training with PyTorch. The main metric of interest is the throughput as it quantifies the training speed but the maximum physical batch size becomes important when training models that are too large to fit even one example at a time. For both metrics DP-SGD becomes more expensive with larger models but the detailed trends differ.
Vision Transformer The throughput difference between Opacus and the non-private baseline with PyTorch (see Figure 2(a)) does not spike but grows steadily as a function of model size, which is interesting considering how big the relative difference is in maximum physical batch size (Figure 3(a)) is: The throughput ranges from a relative difference of for the smallest model to for the largest model while the maximum physical batch size has a relative difference of around for the smallest model and for the largest model.
ResNets We observe a less steady growth trend in terms of throughput with the ResNets in Figure 2(b). When increasing the model size the ResNets do have spikes of growth as the model size grows. The contrast in Figure 2 between ViT and BiT ResNet models is due to the architecture and types of layers. The parameter space grows as the width factor (see Table 1) for the ResNets, so the makes the neural network wider by a factor of three. Based on our results, the width of the layers affects throughput much more than the depth of the network. They have comparable throughput with the same width and different depths, but increasing the width will make the model in the private setting much slower and reduce the maximum batch size significantly.
How much does finding the maximum physical batch size matter? In Figure A.1 in the Appendix, we display the relative throughput as a percentage by dividing the throughput at a particular physical batch size by the maximum achievable throughput. We see that as the physical batch size increases, the throughput will grow as expected, but at some point there is no significant further improvement in throughput from using a larger physical batch size. Practitioners may estimate the optimal batch size based on available memory and performance trade-offs. It is not crucial to set the physical batch size to the maximum possible but a good enough value is fine. Typically, throughput using smaller batches is limited by data loading speeds, but as batch size increases, computation becomes the limiting factor.
4.2 Reasons for the Increase in Computational Cost
Giving a detailed breakdown of low-level operations associated with DP is challenging. However, using GPU profiling tool NVIDIA Nsight System, we can identify three aspects which constitute the majority of DP overheads. Firstly, due to its larger memory footprint, DP-SGD is able to consume smaller physical batches than its non-private counterpart. This results in larger amount of smaller low-level kernel calls which leads to slightly lower utilisation of the GPU. At very small batch sizes even the kernel launch overheads can become a notable factor for slowdown. Secondly, the computation of per-example gradients introduces significant overhead in the backward pass as it cannot be parallelised as in batched gradient computation. This is the most prominent cause of the total overhead. Finally, an additional DP-optimizer step that clips and accumulates the per example gradients, which is not present in the non-DP algorithm, needs to taken after each physical batch (see Table 2).
| Section | PyTorch non-private | Opacus per-example |
| Forward | 81.14 | 101.53 |
| Backward | 163.85 | 681.48 |
| Clip and Accumulation | 0 | 26.76 |
| Optimizer Step | 38.17 | 99.65 |
5 Decreasing the computational cost
This section analyzes the different strategies for training with DP-SGD more efficiently. We evaluate both algorithmic and hardware optimizations and their combinations.
5.1 Optimized gradient clipping algorithms
First, we evaluate the more efficient gradient clipping implementations that have been described in Section 2.2 using the Vision Transformer ViT base model. We chose it as our benchmark model because the middle model size is large enough to evaluate the advantages of the optimized gradient clipping algorithms but does not require excessive amount of time to train. The non-Opacus implementations do not support the BiT ResNet due to their custom weight standardization layer.
| Clipping Mode | V100 | A100 |
| (32GB) | (40GB) | |
| Non private baseline | 216 | 268 |
| Per-example (Opacus) | 28 | 35 |
| Ghost (Private Vision) | 203 | 257 |
| Mix Ghost (Private Vision) | 203 | 257 |
| BK Ghost (FastDP) | 189 | 209 |
| BK Mix Ghost (FastDP) | 189 | 209 |
| BK Mix Opt (FastDP) | 189 | 209 |
Throughput Comparison Figure 4 displays the throughput for each clipping algorithm for each tested GPU. Moving from a V100 to an A100 GPU increased the throughput by times on average over all clipping methods. The one that benefited the most is the per-example clipping by Opacus with a improvement in throughput. This is because of Opacus-specific optimizations. Their implementation is optimized to vectorize the virtual batches and get the most out of the processing unit to compensate for the per-example clipping. Also, we base our virtual batching module on Opacus, which may have further contributed to the advantage seen for Opacus. The other implementations showed benefits similar to those of non-private training. In both GPU architectures, the clipping optimizations consistently maintained their relative throughput difference to their non private baseline. Private Vision gets closer to the non-private baseline physical batch size, but Book Keeping is closer to its throughput with a smaller physical batch size (see Figure 6).
Without sacrificing utility, these optimizations offer an alternative to the original per-example clipping algorithm. Even though Book Keeping performs better, it is by a very narrow margin. Consequently, Private Vision and FastDP remain viable options for implementing ghost clipping. The difference between the two algorithms is the second backward pass over the neural network. Since the Book Keeping trick avoids doing the second backward pass through the network, it has a higher throughput at a small memory cost.
Mixing ghost clipping does not yield any improvement because it determines whether it should apply ghost or per-example clipping, which depends on the size of the inputs and the parameter space. If the dimensions are large enough, the ghost technique will be more expensive (Bu et al., 2022). In a ViT model, the dimensions change less than in a convolutional network. Therefore, despite continually evaluating which method to apply, it always uses ghost clipping. However, if applied to a ResNet model, it should outperform ghost clipping, as it is optimized for convolutional layers. It could not be tested on BiT ResNet models used in this study due to incompatibilities with the Private Vision and FastDP, preventing an assessment of mixed optimization methods.
Maximum physical batch size Table 3 compares the maximum physical batch size for both available GPUs. The maximum physical batch size is larger for the optimizations of DP-SGD than for Opacus because they do not require per-example gradients. Thus, the optimizations allow training much larger models without running out of memory. The maximum physical batch size using Private Vision library is the one that comes closest to the non-private baseline. In general, we can see that the methods are consistent within implementations, with Private Vision and the FastDP reaching the same maximum physical batch size no matter the clipping mode. As expected, the A100 achieves consistently higher maximum physical batch sizes than the V100 due to the larger amount of VRAM.
5.2 Lower precision
We consider using lower precision to speed up computation. We evaluate the use of Tensor Core 32 (TF32) for training. TF32 has 10 bits for precision, with eight range bits, giving it the same range but less precision than 32-bit single precision floats (FP32) (Kharya, 2020). Using lower precision can have benefits exactly where DP training struggles: it requires less memory, uses fewer bits to represent the data, and its operations are optimized for GPU, making them much faster (NVIDIA, 2023). It is specially optimized for the A100 GPU and unavailable for the V100, so we compared training on the A100 with and without TF32.
Experimental results In Figure 5, we display the relative difference between mean throughput between runs with TF32 and FP32. For non-private training, throughput increases with model size. For private training throughput increases for the smaller models, but it goes down again as the model size grows after the base size. Models that are too small do not gain much from TF32, and the larger ones are too expensive. Regarding the memory advantages by TF32, we could not see an improvement. Both models, with and without TF32, could fit the same number of instances.
Concerns regarding TF32 under DP There are two concerns with using lower precision in DP deep learning: its effects on utility and privacy. For the first issue, using lower precision may affect utility, as it is less precise. We did not find a significant decay in the accuracy of the models; it differs by decimal points at the precision (See Table A3 in the Appendix). Regarding privacy, all floating point implementations provide imperfect implementations of real-valued mechanisms, and this can cause additional privacy vulnerabilities (Mironov, 2012). Using lower precision may exacerbate the problem. Discrete mechanisms (e.g. Canonne et al., 2020; Agarwal et al., 2021) that avoid the theoretical challenges exist, but they are often less convenient and may lead to loss of utility, especially in low precision settings. The efficiency of different discrete mechanisms in TF32 is an interesting topic of further research.
6 JAX
In this section we compare the performance of the two JAX implementations with all other DP-SGD frameworks (all of them are based on PyTorch).
Compilation time Comparing JAX to PyTorch requires taking the compilation time into account that the DP-SGD implementations in PyTorch do not utilize. There is no straightforward way of calculating the compilation time, but we measure it as the duration to process the first batch. The execution times for each batch shows that the first one takes much more time than the others, which means it includes the compilation time (Figure A.2). To provide a fair comparison, we also implemented a non-private JAX training using the same virtual batching as PyTorch.
Throughput comparison In Figure 6, both naive JAX implementations (private and non-private) start at the same point, since by using virtual batching, they are processing just one sample at a time and getting the gradients from the compiled function. The non-private benefits more from a larger batch size, although its throughput decreases at 128 instances. The compilation time (see Figure A.2) grows with the batch size. For the private model, it takes more time since the compiled function is more complex than the non-private counterpart. It includes expanding the dimensions and clipping the gradients, while the non-private directly computes the gradient of the whole mini-batch.
The masked DP-SGD jax implementation shows a higher throughput than any other private implementations. Since for this implementation we are fitting the whole logical batch in CPU memory, we can split it and have static sizes. Therefore the compilation time will be higher for the first logical batch, but for the next iterations we can see the gains in speed, as it does not need to recompile. Its throughput changes less with respect to its batch size, in comparison to other methods. It resembles an efficient DP-SGD implementation in JAX using proper Poisson subsampling.
JAX per-example clipping reaches a higher physical batch size, and the throughput is always higher than its Opacus counterpart. However, the Book Keeping Ghost Clipping implementation is closer to the JAX private version. They are the same until the physical batch size of 32, but even then, the variance in the JAX execution is high, meaning that it can even be below the BK-ghost throughput. JAX private naive implementation throughput drops when the batch size doubles to 64. For practitioners that remain in PyTorch, Book Keeping ghost clipping presents a throughput comparable to the execution of the naive JAX while reaching a larger batch size.
Another difference between the two frameworks is the variability in the experiments. PyTorch runs are remarkably consistent in having a low variance, and the same throughput result is expected every time for a fixed seed. JAX executions are more variable than those of PyTorch, likely due to its sensitivity to HPC environment fluctuations and accelerator stochasticity, as noted in Figure A.2. Another contributing factor is JAX’s asynchronous dispatch method, which complicates time benchmarking by issuing a promise rather than immediate results, concealing Python overheads Lastly, the naive method recompiles everytime it process a batch with a different size, adding to the time variability.
Poisson sampling Using JAX for DP introduces complexities, particularly around subsampling which is crucial for privacy accounting. Implementing Poisson subsampling results in variable batch sizes in JAX; changes in batch size require JIT to recompile, leading to graph retracing which is costly and contributes to execution run variability, as discussed by Chua et al. (2024). Our masked DP-SGD implementation overcomes this issue while using proper Poisson subsampling and therefore ensuring the correct privacy budget.
PyTorch Compilation Although compiling PyTorch is possible, we could not see any improvements in terms of speed-up. While compiling the non-private model worked, the speed-up gained was minimal and, in the end, even lower if we consider the compilation time. PyTorch also recompiles after a batch size change. While trying to compile, PyTorch falls back to predefined CUDA operations that are already optimized. In the case of the private setting, the compilation does not recognize Opacus hooks and continues the execution without compiling them.
Leveraging the same kernels to support the private hooks and avoid the compilation would require massive engineering work of writing special kernels for each specific private case. On the other hand, JAX will compile the JIT functions in XLA, but it does not fall back to the kernels, making it more generalizable (Subramani et al., 2021).
7 Multi-GPU Training
This subsection will look at another angle to train deep learning with DP faster: increasing the computational resources enough to decrease the training time. This is relevant when training cost or resource constraints are less important than the time to train a new model.
We utilize V100 GPUs on HPC nodes that have 4 GPUs per node. The other experimental setting is identical to the one in Section 4. Results for utilizing up to 24 A100 GPUs can be found in Figure A.4 in the Appendix. We focus on comparing the scaling behaivour between the non-private baseline that uses PyTorch and the DP-SGD implementation using Opacus. Both frameworks provide mature tooling for distributed training.
Figure 7 shows the throughput increase as a function of number of GPUs. The throughput does not grow linearly and starts changing from the ideal linear scaling after using more than one node (i.e. when using more than 4 GPUs). The communication inside the node is fast, but the communication between nodes will always be slower. The bottleneck is the bandwidth, and it prevents the model from scaling linearly. Notably, it affects the non-private training baseline much more, while the private scales close to optimal up to 32 GPUs. For the 80 GPUs, the private training achieves of the ideal throughput, and the non-private training only achieves . Private training scales better because it is slower and only sometimes saturates the network with updates.
If we use Amdalh’s law to compare the parallelism percentage for each case, we can see that in the private case, we achieve a parallelism compared to a in the non-private case (See Figure A.5 in the Appendix).
8 Conclusion
While DP-SGD is significantly more costly than non-private training, we identified feasible speed-ups that are often easy to apply but have some drawbacks. These are: 1 More efficient implementations of DP-SGD which additionally decrease the memory footprint (allowing for training larger models). However, the implementations are not as mature as Opacus and do not support all types of neural network layers (yet). 2 JAX which processes samples faster than PyTorch but looses the advantage through frequent re-compilations when utilizing proper Poisson sampling, does not offer a comprehensive DP-SGD implementation as PyTorch and exhibits a greater variability in execution times than PyTorch. 3 We present an efficient implementation DP-SGD with JAX that complies with the Poisson sampling requirements while using physical batches of the same length, also complying with JAX efficient optimizations. 4 Lower Precision using TF32 which increases the throughput but the implications on the theoretical guarantees of DP-SGD need to be explored in future work. Finally, we found that distributed computing using DP-SGD scales better than non-private training and allows for fast training of models.
Acknowledgments
This work was supported by the Finnish Ministry of Education and Culture and CSC - IT Centre for Science (Decision diary number OKM/10/524/2022), the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Niki Loppi contributed under the NVIDIA AI Technology Center (NVAITC) Finland program. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. This work has been performed using resources provided by the CSC – IT Center for Science, Finland and the Finnish Computing Competence Infrastructure (FCCI). We thank Talal Alrawajfeh for helpful discussions regarding implementing DP-SGD with JAX.
Appendix A Training Details
A.1 Framework and implementation compatibility
Table A1 shows the corresponding support between implementations and clipping methods. We also include general frameworks, like PyTorch and JAX. Opacus, PrivateVision, and FastDP are PyTorch based implementations.
| Clipping Mode | PyTorch | Opacus | PrivateVision | FastDP | JAX |
| Non-private | |||||
| Per-example | |||||
| Ghost Clipping | |||||
| Mix Ghost | |||||
| Mix Opt |
A.2 Models
Vision Transformer (ViT) (Dosovitskiy et al., 2021). Taken from https://huggingface.co/timm/vit_base_patch16_224.augreg2_in21k_ft_in1k
Big Transfer ResNet (Kolesnikov et al., 2020). Taken from https://github.com/google-research/big_transfer
A.3 Hyperparameters
We use the hyperparameters obtained on request from Tobaben et al. (2023). The hyperparameters for both models are in Table A2. Even though model utility is not the main objective in this work, in the non-private case, the learning rate is suboptimal. By changing it to 0.00027 we see an accuracy improvement, therefore the one we are using.
| Model | Trainable Parameters | Epsilon | Delta | Learning Rate | Max Grad Norm |
| ViT | All | 8 | 0.0003 | 4.63 | |
| ResNet | All | 8 | 0.00098 | 6.53 |
A.4 Poisson Subsampling JAX Algorithm
We present our DP-SGD implementation in JAX that uses the correct Poisson subsampling and therefore we can account for its privacy. The main problem with implementing DP-SGD with JAX is the batches of variable size. In order to address this issue, we compute always full physical batches and mask out gradients so that the total number of used gradients is equal the sampled logical batch sizes. This means that we always compute a little more gradients that required due to sampling. This prevents the recompiling.
Appendix B Additional Results
This section provides additional figures that supplement the findings in the main text.
| Clipping Mode | Test Accuracy |
| Non-private | 0.8050 |
| Non-private JAX | 0.8273 |
| Opacus | 0.8732 |
| Opacus/TF32 | 0.8746 |
| JAX | 0.8009 |
| PV-ghost | 0.8752 |
| BK-ghost | 0.8752 |